OpenStudy (anonymous):

Point M is the midpoint of Ac . Find AC if AM = x+9 and MC = 4x

OpenStudy (anonymous):

M is between A and C. You need to find the distance between A and C. That would be the total distance of the line. You know both sides, just add them together and solve.

OpenStudy (anonymous):

AC is 12

OpenStudy (anonymous):

okay but i keep getting the wrong answer I'm doing x+9+4x and at the end i get 1.8 but thats wrong the right answer is 24 but i don't get that answer

OpenStudy (anonymous):

AC is 12 because midpoint means that M is exactly in the middle so both sides are equal so you equal them together and the answer comes out that AC is 12

OpenStudy (anonymous):

my teacher she gave us a review with the answers and it says the answer is 24

OpenStudy (anonymous):

AC is 24 i just didn't retriceboth together

OpenStudy (anonymous):

I didn't add both sides together*

OpenStudy (anonymous):

The only time you would add them together is if it didn't say M was the midpoint

OpenStudy (anonymous):

could you show me how to do it like type it out cuz I'm not getting those answers

OpenStudy (anonymous):

I missed the midpoint part. AM and MC are the same length. They equal each other. AC would be double the amount of either one. So your equation would be 2(am=ac) or 2(x+9=4x)

OpenStudy (anonymous):

Solving x+9=4x would give you the length of either side. The 2 would double it to the find the total length.

OpenStudy (anonymous):

x+9=4x then subtract x from both sides to get 9=3x then divide 3 on both sides then x equals 3 then you do AM= (3)+9= 12 MC=4(3)= 12 then add them both together to get AC=24

OpenStudy (anonymous):

thank you both so much!!!

OpenStudy (anonymous):

You're Welcome I'm taking Geometry now and this is exactly what we are doing

OpenStudy (anonymous):

could you help me with 2 more problems ?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

point J is between H and K on HK . If HJ= 5x-4, jk= 8x-10, and Kh = 38 find jk

OpenStudy (anonymous):

Point B is between A and C on AC. If ac= 4x+4, ab= 14 and bc = 3x-4 find ac

OpenStudy (anonymous):

1. 5x-4+8x-10=38, combine like terms 13x-14=38 add 14 to both sides 13x=52 divide 13 on both sides and x=4. JK=8(4)-10= 22

OpenStudy (anonymous):

Start by drawing it out. Might help you see it better.

OpenStudy (anonymous):

2. 14+3x-4=4x+4 combine like terms 3x+10=4x+4 subtract 3x from both sides to get 10=x+4 subtrat 4 fom both sides to get x=6 then plug it in to AC, AC=4(6)+4=28

OpenStudy (anonymous):

thnx soo much! i need so much help on math tbh

OpenStudy (anonymous):

The way I think about these is to draw it out. Then I can see that HJ+JK=HK. Put in the equations as they are given and solve. With that you can solve for x, and then just plug x into the equation they are asking you to solve (in that case jk). I'm a very visual person and find it much easier that way.

OpenStudy (anonymous):

You're welcome im really good at all math, Algebra, Geometry and Trig. Im only in 10th grade too. Im also a visual learner and im sitting here writing these on paper lol

OpenStudy (anonymous):

lol! see i don't get it at all… like find the coordinates of the midpoint of PQ if p (-5,5) and Q (7,3) help

OpenStudy (anonymous):

Lol do you have a reference sheet? theres a formula for midpoint when using coordinates

OpenStudy (anonymous):

I can answer that if you want me to.

OpenStudy (anonymous):

-5+7 5+3 ------ ------- 2 , 2 (1,4)

OpenStudy (anonymous):

how did u get (1,4) though

OpenStudy (anonymous):

you add -5+7 then divide that by 2 to get 1. Then 5+3 to get 8 and divide that by 2 to get 4

OpenStudy (anonymous):

It's easier to understand if you had a refereance sheet to look at all the time

OpenStudy (anonymous):

thanks I'm going to fail my exam 2maro.. are u willing to keep helping me

OpenStudy (anonymous):

Ill help you for as long as you need. All I have is time since we don't have school today, and I took all my midterms last week what grade are you in?

OpenStudy (anonymous):

I'm in 9th, lol and thanks so much! all our exams are this week ! and I've been studying for math all weekend and i just can't get it..

OpenStudy (anonymous):

Im sure you'll get it sooner or later

OpenStudy (anonymous):

hahah hopefully ! what is the distance between F(-1,4) and G(6,-4) on the circle wtf is this omg help.

OpenStudy (anonymous):

it also says round to nearest hundredth

OpenStudy (anonymous):

On a circle it's kind of like the radius but the midpoint is still a point.

OpenStudy (anonymous):

so do u solve it with the (x1,y1) thing

OpenStudy (anonymous):

No thats to find the midpoint for this question you use the distance formula

OpenStudy (anonymous):

ohhh okay i got it!

OpenStudy (anonymous):

do you still need me to help you solve it?

OpenStudy (anonymous):

umm yeah please

OpenStudy (anonymous):

√ ̅(6+1)^2 +(-4-4)^2 plug it in to the calculator and the answer comes out to be 12.81

OpenStudy (anonymous):

omg ur a life savior

OpenStudy (anonymous):

Hahah thanks

OpenStudy (anonymous):

so for this we have to find x and at the top of the triangle theres (x) then theres 2x and its a right triangle

OpenStudy (anonymous):

nvm i got it lol

OpenStudy (anonymous):

Like this

OpenStudy (anonymous):

Oh OK

OpenStudy (anonymous):

If RG is the median of triangle TRI , TG=(5-10X), and TI=(32-9X) FIND X

OpenStudy (anonymous):

.. sorry for caps lmao

OpenStudy (anonymous):

is there a picture?

OpenStudy (anonymous):

ya one sec

OpenStudy (anonymous):

its not letting me draw it

OpenStudy (anonymous):

Ill draw waht I think it looks like

OpenStudy (anonymous):

Thats what I thought it looked like.

OpenStudy (anonymous):

haha yeah so could u help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

X=-2. 5-10x+5-10x=32-9x combine like terms, -20x+10=32-9x add 9x to both sides -11+10=32 subtract 10 to both sides to get -11=22 then divide -11 to both sides to get x=-2.

OpenStudy (anonymous):

thank you I'm done with math lol thnx!